# Towards deterministic non-Gaussianity on a chip

### v1

@@ Acknowledgements.
S.E.F. and J.E.S. acknowledge the Natural Sciences and Engineering Research Council of Canada European Union’s Horizon Europe Research and Innovation Programme (101070700, project MIRAQLS) for financial support, and thank Colin Vendromin for insightful discussions. S.E.F. acknowledges support from a Walter C. Sumner Memorial Fellowship. M.L. acknowledges support by PNRR MUR Project No. PE0000023-NQSTI. M.B. acknowledges support from the Quantum Research and Development Initiative, led by the National Research Council Canada, under the National Quantum Strategy, and thanks Nicolás Quesada and Boris Braverman for stimulating discussions.
Appendix B: Nonlinear physics— The relevant Hamiltonians for the cascaded SFWM processes are
@@
$H_{\text{SFWM1}}$ $=-\frac{1}{4\epsilon_{0}}\frac{4!}{2!1!1!}\int\mathrm{d}\bm{r}\Gamma^{(3)}_{ijkl}(\bm{r})D^{\text{out},i}_{S_{1}}(\bm{r})D^{\text{out},j}_{I}(\bm{r})$
$\hskip 85.35826pt\times D^{\text{in},k}_{P_{1}}(\bm{r})D^{\text{in},l}_{P_{1}}(\bm{r}),$ (18)
@@
$H_{\text{SFWM2}}$ $=-\frac{1}{4\epsilon_{0}}\frac{4!}{1!1!1!1!}\int\mathrm{d}\bm{r}\Gamma^{(3)}_{ijkl}(\bm{r})D^{\text{out},i}_{S_{2}}(\bm{r})D^{\text{out},j}_{S_{3}}(\bm{r})$
$\hskip 85.35826pt\times D^{\text{out},k}_{I}(\bm{r})D^{\text{in},l}_{P_{2}}(\bm{r}).$ (19)
@@
$\bm{D}_{u}^{\text{in/out}(I)}(\bm{r},t)=$ $\sum_{\lambda}\int\mathrm{d}k\bm{\mathcal{D}}_{uk}^{\text{in/out},\lambda}(\bm{r})$
@@
$\bm{D}_{u}^{\text{in}(I)}(\bm{r},t)=\alpha_{u}\int\mathrm{d}k\bm{\mathcal{D}}_{uk}^{\text{in},\lambda}(\bm{r})\phi_{u}(k)e^{-i\omega_{uk}t}+\mathrm{H.c..}$ (22)
@@
$H_{\text{SFWM1}}^{(I)}$ $(t)=\frac{-\alpha_{P_{1}}^{2}}{3\epsilon_{0}}\int\mathrm{d}\bm{r}\Gamma_{ijkl}^{(3)}(\bm{r})$
$\times$ $\left(\sum_{\lambda}\int\mathrm{d}k{\mathcal{D}}_{S_{1}k}^{\text{out},\lambda,i}(\bm{r})a_{S_{1}}^{\text{out},\lambda}(k)e^{-i\omega_{S_{1}k}t}+\mathrm{H.c.}\right)$
$\times$ $\left(\sum_{\lambda^{\prime}}\int\mathrm{d}k{\mathcal{D}}_{Ik}^{\text{out},\lambda^{\prime},j}(\bm{r})a_{I}^{\text{out},\lambda^{\prime}}(k)e^{-i\omega_{Ik}t}+\mathrm{H.c.}\right)$
$\times$ $\left(\int\mathrm{d}k{\mathcal{D}}_{P_{1}k}^{\text{in},k}(\bm{r})\phi_{P_{1}}^{\text{in}}(k)e^{-i\omega_{P_{1}k}t}+\mathrm{H.c.}\right)$
$\times$ $\left(\int\mathrm{d}k{\mathcal{D}}_{P_{1}k}^{\text{in},l}(\bm{r})\phi_{P_{1}}^{\text{in}}(k)e^{-i\omega_{P_{1}k}t}+\mathrm{H.c.}\right).$ (23)
@@
$K_{1}^{\lambda\lambda^{\prime}}(k_{1},...,k_{4})=\frac{3}{\epsilon_{0}}\int\mathrm{d}\bm{r}$ $\Gamma_{ijkl}^{(3)}\left[\mathcal{D}^{\text{out},\lambda,i}_{S_{1}k_{1}}\mathcal{D}^{\text{out},\lambda^{\prime},j}_{Ik_{2}}\right]^{*}$
@@
$K_{2}^{\mu\mu^{\prime}\mu^{\prime\prime}}(k_{1},...,k_{4})=\frac{6}{\epsilon_{0}}\int\mathrm{d}\bm{r}$ $\Gamma_{ijkl}^{(3)}\left[\mathcal{D}^{\text{out},\mu^{\prime},i}_{S_{2}k_{1}}\mathcal{D}^{\text{out},\mu^{\prime\prime},j}_{S_{3}k_{2}}\right]^{*}$
@@
We neglect the final term in Eq. (30), which describes the generation of two pairs in the first ring. Inserting Eqs. (24) and (25) into (30) and taking $_{0}\rightarrow-\infty, tt\rightarrow\infty$, we obtain Eq. (4), with
@@
Appendix C Supplemental material 1: Asymptotic fields for dual rings coupled to a waveguide
$\bm{D}(\bm{r})=\sum_{u}\bm{D}_{u}^{\mathrm{in/out}}(\bm{r})+\mathrm{H.c.},$ (35)
with
$\bm{D}_{u}^{\mathrm{in/out}}(\bm{r})=\sum_{\lambda}\int\mathrm{d}k\bm{\mathcal{D}}_{uk}^{\mathrm{in/out},\lambda}(\bm{r})a_{u}^{\mathrm{in/out},\lambda}(k)+\mathrm{H.c..}$ (36)
We provide the fields inside the microring resonators for each case in Fig. 4, obtained from the Heisenberg equations for the specific asymptotic state [27]. Using $\bm{r}_{\perp}$ to denote a vector along the cross-section of the waveguide constructing the ring, and $\zeta$ to identify the azimuthal coordinate along the circumference of the ring (see Fig. 1a), the asymptotic-in fields for the actual waveguide (Fig. 4a) are
$\bm{\mathcal{D}}_{uk}^{\mathrm{in,ac}}(\bm{r})=-\sqrt{\frac{\hbar\omega_{u}}{4\pi}}\bm{\mathsf{d}}^{(1)}_{u}(\bm{r}_{\perp})F_{u-}^{\mathrm{ac}}(k)e^{-i(k-K_{u})a}e^{i\kappa^{(1)}_{u}\zeta},$
$[\bm{r}\in\,\mathrm{ring\,1}]$ (37)
$\bm{\mathcal{D}}_{uk}^{\mathrm{in,ac}}(\bm{r})=-\sqrt{\frac{\hbar\omega_{u}}{4\pi}}\bm{\mathsf{d}}^{(2)}_{u}(\bm{r}_{\perp})F_{u-}^{\mathrm{ac}(2)}(k)\left(1+\frac{i\gamma_{u}^{\mathrm{ac}(1)}}{v_{u}^{\mathrm{ac}}}\sqrt{\mathcal{L}^{(1)}}F_{u-}^{\mathrm{ac}(1)}(k)\right)e^{i(k-K_{u})a}e^{i\kappa^{(2)}_{u}\zeta},$
$[\bm{r}\in\,\mathrm{ring\,2}]$ (38)
the asymptotic-out fields for the actual waveguide (Fig. 4b) are
$\bm{\mathcal{D}}_{uk}^{\mathrm{out,ac}}(\bm{r})=-\sqrt{\frac{\hbar\omega_{u}}{4\pi}}\bm{\mathsf{d}}^{(1)}_{u}(\bm{r}_{\perp})F_{u+}^{\mathrm{ac}(1)}(k)\left(1-\frac{i\gamma_{u}^{\mathrm{ac}(2)}}{v_{u}^{\mathrm{ac}}}\sqrt{\mathcal{L}^{(2)}}F_{u+}^{\mathrm{ac}(2)}(k)\right)e^{-i(k-K_{u})a}e^{i\kappa^{(1)}_{u}\zeta},$
$[\bm{r}\in\,\mathrm{ring\,1}]$ (39)
$\bm{\mathcal{D}}_{uk}^{\mathrm{out,ac}}(\bm{r})=-\sqrt{\frac{\hbar\omega_{u}}{4\pi}}\bm{\mathsf{d}}^{(2)}_{u}(\bm{r}_{\perp})F_{u+}^{\mathrm{ac}(2)}(k)e^{i(k-K_{u})a}e^{i\kappa^{(2)}_{u}\zeta},$
$[\bm{r}\in\,\mathrm{ring\,2}]$ (40)
$\bm{\mathcal{D}}_{uk}^{\mathrm{out,ph}(1)}(\bm{r})=-\sqrt{\frac{\hbar\omega_{u}}{4\pi}}\bm{\mathsf{d}}^{(1)}_{u}(\bm{r}_{\perp})F_{u+}^{\mathrm{ph}(1)}(k)e^{-i(k-K_{u})a}e^{i\kappa^{(1)}_{u}\zeta},$
$[\bm{r}\in\,\mathrm{ring\,1}]$ (41)
$\bm{\mathcal{D}}_{uk}^{\mathrm{out,ph}(1)}(\bm{r}_{\perp})=0,$
$[\bm{r}\in\,\mathrm{ring\,2}]$ (42)
$\bm{\mathcal{D}}_{uk}^{\mathrm{out,ph}(2)}(\bm{r})=-\sqrt{\frac{\hbar\omega_{u}}{4\pi}}\bm{\mathsf{d}}^{(1)}_{u}(\bm{r}_{\perp})F_{u+}^{\mathrm{ac}(1)}(k)\left(-\frac{i\gamma_{u}^{\mathrm{ac}(2)}}{v_{u}^{\mathrm{ac}}}\sqrt{\mathcal{L}^{(2)}}F_{u+}^{\mathrm{ph}(2)}(k)\right)e^{-i(k-K_{u})a}e^{i\kappa^{(1)}_{u}\zeta}$
$[\bm{r}\in\,\mathrm{ring\,1}]$ (43)
$\bm{\mathcal{D}}_{uk}^{\mathrm{out,ph}(2)}(\bm{r})=-\sqrt{\frac{\hbar\omega_{u}}{4\pi}}\bm{\mathsf{d}}^{(2)}_{u}(\bm{r}_{\perp})F_{u+}^{\mathrm{ph}(2)}(k)e^{i(k-K_{u})a}e^{i\kappa^{(2)}_{u}\zeta},$
$[\bm{r}\in\,\mathrm{ring\,2}]$ (44)
where $\bm{\mathsf{d}}^{(\ell)}_{u}(\bm{r})$ are the spatial mode profiles in the $\ell^{\mathrm{th}}$ ring ($\ell\in\{1,2\}$) associated to the frequency range $u$ [24]. We have also introduced a field enhancement factor
The spatial modes are normalized according to
$\begin{split}\int\frac{v_{p}(\bm{r}_{\perp};\omega_{u})}{v_{g}(\bm{r}_{\perp};\omega_{u})}\frac{\bm{\mathsf{d}}_{u}^{(\ell)*}(\bm{r}_{\perp})\cdot\bm{\mathsf{d}}_{u}^{(\ell)}(\bm{r}_{\perp})}{\epsilon_{0}\varepsilon_{1}(\bm{r}_{\perp};\omega_{u})}\mathrm{d}\bm{r}_{\perp}=1,\end{split}$ (46)
Appendix D Supplemental material 2: Tensor and nonlinear parameter
The III-V semiconductors of interest such as AlGaAs and InGaP belong to the $\bar{4}3$m point group and have a cubic symmetry [51]. In general, the $\chi^{(3)}$ tensor has 21 nonzero components, of which only one component is independent assuming Kleinman symmetry. We have $\chi^{(3)}_{x^{\prime}x^{\prime}x^{\prime}x^{\prime}}=\chi^{(3)}_{y^{\prime}y^{\prime}y^{\prime}y^{\prime}}=\chi^{(3)}_{z^{\prime}z^{\prime}z^{\prime}z^{\prime}}=3\chi^{(3)}_{[x^{\prime}x^{\prime}y^{\prime}y^{\prime}]}=3\chi^{(3)}_{[x^{\prime}x^{\prime}z^{\prime}z^{\prime}]}=3\chi^{(3)}_{[y^{\prime}y^{\prime}z^{\prime}z^{\prime}]}=\bar{\chi}^{(3)}$, where the notation $[abcd]$ denotes the distinct permutations of the set $\{a,b,c,d\}$.
We can introduce a laboratory frame that coincides with the crystal’s axes $x^{\prime}y^{\prime}z^{\prime}$, and consider a crystal grown in the $z^{\prime}$ direction. We also introduce a ring frame xyzxyz where the $zx$ plane forms the cross-section of the waveguide constructing the ring, and $y$ is in the direction of the azimuthal angle going around the ring’s circumference (see Fig. 1). Since we assume Kleinman symmetry with a cubic crystal, the $\chi^{(3)}$ tensor takes a form similar to one of an isotropic material, and is therefore invariant under rotations of the coordinate system (e.g., about the $z^{\prime}$-axis). Consequently, in our case with the defined coordinate systems, the transformed tensor components satisfy $\chi^{(3)}_{i^{\prime}j^{\prime}k^{\prime}l^{\prime}}=\chi^{(3)}_{ijkl}$.
We can write Eqs. (28) and (29) in terms of nonlinear parameters $\gamma_{\mathrm{NL}}$ [24]. For simplicity, we can write the fields from Eqs. (37)–(44) as
$\bm{\mathcal{D}}_{uk}^{\mathrm{in/out},\lambda}(\bm{r})=\sqrt{\frac{\hbar\omega_{u}}{4\pi}}\bm{\mathsf{d}}^{(\ell)}_{u}(\bm{r})\tilde{{\mathcal{D}}}_{uk}^{\mathrm{in/out},\lambda}e^{i\kappa^{(\ell)}_{u}\zeta}\,\,\,\,[\bm{r}\in\,\mathrm{ring}\,\ell],$ (47)

### v2

@@ Acknowledgements.
S.E.F. and J.E.S. acknowledge the Natural Sciences and Engineering Research Council of Canada, and the European Union’s Horizon Europe Research and Innovation Programme (101070700, project MIRAQLS) for financial support, and thank Colin Vendromin for insightful discussions. S.E.F. acknowledges support from a Walter C. Sumner Memorial Fellowship. M.L. acknowledges support by PNRR MUR Project No. PE0000023-NQSTI. M.B. acknowledges support from the Quantum Research and Development Initiative, led by the National Research Council Canada, under the National Quantum Strategy, and thanks Nicolás Quesada and Boris Braverman for stimulating discussions.
Appendix A Linear physics
The linear Hamiltonian for the rings, point-coupled to a waveguide as shown in Fig. 1a, consists of the sum of the isolated ring Hamiltonians, three waveguide Hamiltonians, two coupling Hamiltonians to the bus waveguide, and two coupling Hamiltonians to the phantom waveguides ($\lambda\in\{\mathrm{ac,ph1,ph2}\}$, $\ell\in\{1,2\}$):
[…]
The $\ell^{\mathrm{th}}$ isolated ring Hamiltonian is
$\begin{split}&H^{(\ell)}_{\mathrm{ring}}=\sum_{u}\hbar\omega_{u}^{(\ell)}b^{(\ell)\dagger}_{u}b_{u}^{(\ell)},\end{split}$ (8)
with the $\ell^{\mathrm{th}}$ ring’s resonant frequencies denoted by $\omega_{u}^{(\ell)}$, with the usual commutation relations
$\left[b_{u}^{(\ell)},b_{u^{\prime}}^{(\ell^{\prime})\dagger}\right]=\delta_{uu^{\prime}}\delta_{\ell\ell^{\prime}}.$ (9)
The index $u$ denotes a frequency range with center resonance $\omega_{u}$ where up to single a resonance of each ring can be found; e.g., in the frequency range $S_{1}$, the first ring could have a resonance with resonant frequency $\omega_{S_{1}}^{(1)}$ or not, and similarly for $\omega_{S_{1}}^{(2)}$, and in general both frequencies can be different.
The Hamiltonian for the isolated $\lambda$ waveguide is
$\begin{split}&H_{\lambda}^{\mathrm{wg}}=\sum_{u}\Bigg[\int\hbar\omega_{u}\psi^{\lambda\dagger}_{u}(s)\psi^{\lambda}_{u}(s)\mathrm{d}s\\ &-\frac{i\hbar v_{u}}{2}\int\left(\psi^{\lambda\dagger}_{u}(s)\frac{\partial\psi^{\lambda}_{u}(s)}{\partial s}-\frac{\partial\psi_{u}^{\lambda\dagger}(s)}{\partial s}\psi^{\lambda}_{u}(s)\right)\mathrm{d}s\Bigg],\end{split}$ (10)
where
$\psi^{\lambda}_{u}(s)=\int\frac{\mathrm{d}k}{\sqrt{2\pi}}a^{\lambda}_{u}(k)e^{i(k-K_{u})s},$ (11)
is the channel operator associated with the resonant frequency range of the resonator [39], and has the usual commutation relations
$\left[a_{u}^{\lambda}(k),\left(a_{u^{\prime}}^{\lambda^{\prime}}(k)\right)^{\dagger}\right]=\delta_{uu^{\prime}}\delta_{\lambda\lambda^{\prime}}\delta(k-k^{\prime}).$ (12)
We introduced $v_{u}=\left(\partial\omega_{uk}/\partial k\right)_{K_{u}}$ and $K_{u}$, which are the group velocity and a reference wavenumber in the waveguides associated with the frequency range $u$, over which group velocity dispersion is neglected. For convenience, $K_{u}$ is often chosen as a resonant wavenumber of a ring $\kappa_{u}^{(\ell)}=2\pi m_{u}^{(\ell)}/\mathcal{L}^{(\ell)}$, where $m_{u}^{(\ell)}$ and $\mathcal{L}^{(\ell)}$ are the resonant mode numbers and circumference of the $\ell^{\mathrm{th}}$ ring, respectively. The waveguides are assumed to be identical, and so there is no waveguide dependence on $v_{u}$ and $K_{u}$, but this can easily be generalized. We write the dispersion relation around the frequency $\omega_{u}$ of the frequency range of interest
$\omega_{uk}=\omega_{u}+v_{u}(k-K_{u}).$ (13)
Using $s$ to denote the distance along each waveguide (bus waveguide and phantom waveguides), for the bus waveguide we take the coupling point to the first (second) ring to be at s=as=-a ($s=+a$), and for both phantom waveguides the coupling point is at $s=0$. The coupling Hamiltonians for the $\ell^{\mathrm{th}}$ ring are
[…]
where we take a-a (+a+a) for $\ell=1$ ($\ell=2$). We define decay rates to the waveguides, quality factors, and escape efficiencies from the coupling coefficients $\gamma_{u}^{\lambda(\ell)}$ from the above Hamiltonians [39]:
$\Gamma_{u}^{\lambda(\ell)}$ $\equiv\frac{|\gamma_{u}^{\lambda(\ell)}|^{2}}{2v_{u}}=\frac{\omega_{u}^{(\ell)}}{2Q_{u}^{\lambda(\ell)}},$ (15)
$\bar{\Gamma}^{(\ell)}_{u}$ $\equiv\Gamma_{u}^{\mathrm{ac}(\ell)}+\Gamma_{u}^{\mathrm{ph}(\ell)}=\frac{\omega_{u}^{(\ell)}}{2Q_{u}^{\mathrm{(load)}{(\ell)}}},$ (16)
$\eta_{u}^{\lambda(\ell)}$ $\equiv\frac{\Gamma_{u}^{\lambda(\ell)}}{\bar{\Gamma}_{u}^{(\ell)}}=\frac{Q_{u}^{\mathrm{(load)}{(\ell)}}}{Q_{u}^{\lambda(\ell)}}.$ (17)
Appendix B Nonlinear physics
The relevant Hamiltonians for the cascaded SFWM processes are
@@ Appendix B Nonlinear physics
$H_{\text{SFWM1}}$ $=-\frac{1}{4\epsilon_{0}}\frac{4!}{2!1!1!}\int\mathrm{d}\boldsymbol{r}\Gamma^{(3)}_{ijkl}(\boldsymbol{r})D^{\text{out},i}_{S_{1}}(\boldsymbol{r})D^{\text{out},j}_{I}(\boldsymbol{r})$
$\hskip 85.35826pt\times D^{\text{in},k}_{P_{1}}(\boldsymbol{r})D^{\text{in},l}_{P_{1}}(\boldsymbol{r}),$ (18)
@@
$H_{\text{SFWM2}}$ $=-\frac{1}{4\epsilon_{0}}\frac{4!}{1!1!1!1!}\int\mathrm{d}\boldsymbol{r}\Gamma^{(3)}_{ijkl}(\boldsymbol{r})D^{\text{out},i}_{S_{2}}(\boldsymbol{r})D^{\text{out},j}_{S_{3}}(\boldsymbol{r})$
$\hskip 85.35826pt\times D^{\text{out},k}_{I}(\boldsymbol{r})D^{\text{in},l}_{P_{2}}(\boldsymbol{r}).$ (19)
We note that in deriving Eqs. (18)-(19) full permutation symmetry of the $\Gamma^{(3)}$ tensor is assumed [39], and in Appendix C we briefly discuss the effects of this assumption on our results.
@@
$\boldsymbol{D}_{u}^{\text{in/out}(I)}(\boldsymbol{r},t)=$ $\sum_{\lambda}\int\mathrm{d}k\boldsymbol{\mathcal{D}}_{uk}^{\text{in/out},\lambda}(\boldsymbol{r})$
@@
$\boldsymbol{D}_{u}^{\text{in}(I)}(\boldsymbol{r},t)=\alpha_{u}\int\mathrm{d}k\boldsymbol{\mathcal{D}}_{uk}^{\text{in},\lambda}(\boldsymbol{r})\phi_{u}(k)e^{-i\omega_{uk}t}+\mathrm{H.c..}$ (22)
@@
$H_{\text{SFWM1}}^{(I)}$ $(t)=\frac{-\alpha_{P_{1}}^{2}}{3\epsilon_{0}}\int\mathrm{d}\boldsymbol{r}\Gamma_{ijkl}^{(3)}(\boldsymbol{r})$
$\times$ $\left(\sum_{\lambda}\int\mathrm{d}k{\mathcal{D}}_{S_{1}k}^{\text{out},\lambda,i}(\boldsymbol{r})a_{S_{1}}^{\text{out},\lambda}(k)e^{-i\omega_{S_{1}k}t}+\mathrm{H.c.}\right)$
$\times$ $\left(\sum_{\lambda^{\prime}}\int\mathrm{d}k{\mathcal{D}}_{Ik}^{\text{out},\lambda^{\prime},j}(\boldsymbol{r})a_{I}^{\text{out},\lambda^{\prime}}(k)e^{-i\omega_{Ik}t}+\mathrm{H.c.}\right)$
$\times$ $\left(\int\mathrm{d}k{\mathcal{D}}_{P_{1}k}^{\text{in},k}(\boldsymbol{r})\phi_{P_{1}}^{\text{in}}(k)e^{-i\omega_{P_{1}k}t}+\mathrm{H.c.}\right)$
$\times$ $\left(\int\mathrm{d}k{\mathcal{D}}_{P_{1}k}^{\text{in},l}(\boldsymbol{r})\phi_{P_{1}}^{\text{in}}(k)e^{-i\omega_{P_{1}k}t}+\mathrm{H.c.}\right).$ (23)
@@
$K_{1}^{\lambda\lambda^{\prime}}(k_{1},...,k_{4})=\frac{3}{\epsilon_{0}}\int\mathrm{d}\boldsymbol{r}$ $\Gamma_{ijkl}^{(3)}\left[\mathcal{D}^{\text{out},\lambda,i}_{S_{1}k_{1}}\mathcal{D}^{\text{out},\lambda^{\prime},j}_{Ik_{2}}\right]^{*}$
@@
$K_{2}^{\mu\mu^{\prime}\mu^{\prime\prime}}(k_{1},...,k_{4})=\frac{6}{\epsilon_{0}}\int\mathrm{d}\boldsymbol{r}$ $\Gamma_{ijkl}^{(3)}\left[\mathcal{D}^{\text{out},\mu^{\prime},i}_{S_{2}k_{1}}\mathcal{D}^{\text{out},\mu^{\prime\prime},j}_{S_{3}k_{2}}\right]^{*}$
@@
We neglect the final term in Eq. (30), which describes the generation of two pairs in the first ring. Inserting Eqs. (24) and (25) into (30) and taking $t_{0}\rightarrow$-$\infty$, we obtain Eq. (4), with
@@
Appendix C Nonlinear parameters and tensors
We can write Eqs. (28) and (29) in terms of nonlinear parameters $\gamma_{\mathrm{NL}}$ [39]. For simplicity, we write the fields from Eqs. (60)–(67) as
$\boldsymbol{\mathcal{D}}_{uk}^{\mathrm{in/out},\lambda}(\boldsymbol{r})=\sqrt{\frac{\hbar\omega_{u}}{4\pi}}\boldsymbol{\mathsf{d}}^{(\ell)}_{u}(\boldsymbol{r})\tilde{{\mathcal{D}}}_{uk}^{\mathrm{in/out},\lambda}e^{i\kappa^{(\ell)}_{u}\zeta}\,\,\,\,[\boldsymbol{r}\in\,\mathrm{ring}\,\ell],$ (35)